# Navigating the Future of Federated Recommendation Systems with Foundation Models

### v2

@@ III-B1 Taxonomy of FM
As shown in Fig. 6, similar to FRS, FM can also be classified based on data type or functionality. Specifically, based on the type of data used during training, FM is mainly divided into followings [18]:
  • Language FM: These types of FM mainly deal with textual data and are trained on large textual datasets to understand and generate language [19, 20]. They are very effective in natural language processing (NLP) tasks such as machine translation [62], search [63, 64], and sentiment analysis [65]. Typical language models include BERT [19], GPT-3 [56], and LLama [66], which typically use deep transformer architectures that are able to capture complex linguistic regularities and demonstrate excellent performance on multilingual tasks.
  • Vision FM: Vision FM focuses on processing and understanding image data. These models are able to perform tasks such as object recognition [21, 67, 68], image segmentation [69, 70, 71], and visual reasoning [72, 73] by learning large amounts of image data. For example, DINOv2 [74] and SAM [69] are basic models trained specifically for visual tasks, and using self-supervised learning methods, these models learn valid visual representations without labelling the data.
  • Multi-modal FM: Multi-modal models can simultaneously process and understand multiple types of data, such as text and images. Such models can excel in cross-modal tasks such as image captioning and visual quizzing [75, 76, 77] by integrating information from different modalities. CLIP [23] and DALL·E [24] are representative of such models that are able to understand the relationship between an image and its corresponding textual descriptions, demonstrating flexibility and robustness in handling multiple data types.
Functionally, FMs are generally categorized into two types [18, 78, 38]:
  • Discriminative FM: The main task of discriminative FM is to distinguish or predict specific outputs from given input data. These models are generally based on the BERT family and are more concerned with learning decision boundaries from the data to perform tasks like classification [79, 80] and regression [81].
  • Generative FM: The core goal of generative FM is to learn the distribution of the data and be able to generate new data samples [25]. These models, such as GPT-3, DALL·E, etc., are able to capture the underlying structure of the data and thus generate new instances that are similar but different from the training data. Generative FM has a wide range of applications in areas such as the generation of text [82] and images [12, 13].
@@ III-B2 Adaptation
Fig. 1 illustrates a commonly used technique for pre-trained models, known as the Adaptation. This technique adds new lightweight layers while maintaining the original parameters of the pre-trained model, enabling fine-tuning and extension for specific tasks [83, 41]. This approach is suitable for multi-task learning and tasks performed in resource-limited environments, such as few-shot learning [84]. Adaptations have found wide applications in natural language processing and computer vision tasks [85, 86, 87] . According to [84], adaptation methods are categorized into three main types:
  • Prompt-based Fine-tuning Adaptation Methods are inspired by the principles of prompt learning in NLP, adapts the model to a specific downstream task by using a portion of the fixed text input as a learnable vector and fine-tuning it using downstream task data. For instance, CoOp [88] introduces the concept of prompt learning into the adaptation of downstream tasks for multimodal pre-trained foundation models by automatically constructing contextual prompts using learnable word embeddings. CoCoOp [89] builds upon this by creating a meta-network that learns image features, which are then combined with prompt vectors to enhance the generalization performance of CoOp on new category data.
  • Adapter-based Fine-tuning Adaptation Methods entail the incorporation of small, trainable modules called Adapters into a pre-trained model. These adapters are subjected to fine-tuning, rather than the entire model, thereby enabling the model to adapt to novel tasks with minimal alterations to its original parameters. CLIP-Adapter [87] adds simple adapter structures within the CLIP model that are fine-tuned using a small dataset. This method allows the model to adapt to new tasks by learning task-specific features without the need for extensive retraining of the entire model. CALIP [90] enhances the cross-modal interaction capabilities of adapters by integrating text and image features using attention maps and incorporating two trainable linear layers before and after this integration. LoRA [85] adjusts the weights of a pre-trained model by multiplying the weight matrices with a low-rank matrix, allowing for fine-tuning of the model without significantly increasing the number of parameters.
  • External Knowledge-based Adaptation Methods integrate external knowledge (e.g., from knowledge graphs or additional language model queries) into the adaptation process. This is particularly advantageous for enhancing the model’s comprehension and efficacy in tasks where domain-specific knowledge is vital but data is scarce. CuPL [91]enhances zero-shot classification performance by asking questions to a LLM and generating multiple descriptive statements for each category, enriching the semantics of the categories. LaBo [92] uses GPT-3 to generate a large pool of candidate feature descriptors. It then selects the optimal subset of descriptors for classification decisions, considering the distinctiveness of features for other categories and the coverage for the current category.
@@ IV Federated Recommendation Systems with Foundation Models
As described earlier in Fig. 2, a typical FRS typically consists of three stages [27]: client model update, communication, and global aggregation. Integration with FM should also occur at these three stages.
@@ IV-A Client Model Update
  • As mentioned earlier, a major advantage of FRS is privacy. The data required for model training in FRS, such as user information and interaction history, is privacy sensitive and therefore must remain on the user’s device, creating data silos. This approach not only complies with data privacy regulations (such as GDPR), but also increases user trust. However, this means that each user has a relatively small amount of data compared to the entire dataset, and each user only accesses a portion of the item set, leading to data sparsity.
@@
  • Each client independently trains the model on its local data. This allows the model to better adapt to each user’s unique preferences and behaviors, improving personalization and accuracy of recommendations. The resulting data distribution often does not satisfy the independent and identically distributed (IID) assumption, leading to the problem of data heterogeneity.
@@
  • In addition, user devices are typically consumer products such as mobile phones and personal computers, which are characterized by unstable communication and limited computational resources. This requires keeping the computational load of the models deployed on the client and the amount of information exchanged with the server as low as possible.
@@
FM is typically pre-trained on large, diverse datasets, providing it with rich knowledge and strong representational capabilities. This pre-training provides it with prior knowledge that allows it to quickly adapt to specific customer data through fine-tuning [18]. As a result, FM can learn general feature representations that provide a degree of adaptability to different data distributions. Applying these models during the client update phase in FRS can significantly improve the performance of local models. Specifically:
@@
  • FM can apply knowledge learned from large corpora to local data through transfer learning. For example, a pre-trained FM can be effectively fine-tuned on a small amount of local data in the client to adapt to specific user behavior data, thereby achieving good performance in specific downstream tasks such as providing more accurate recommendations [38]. In addition, by fine-tuning the base model locally, sensitive data does not need to leave the device, protecting user privacy [93]. As a result, customers can achieve better model performance with limited data.
@@
  • FM can also capture complex user preferences and behaviors. For example, FM has strong semantic understanding capabilities that allow them to better understand users’ search queries [63, 64], comments [37], and other textual data [94] to provide recommendations that better meet user needs. In addition, as user behavior changes, FM can quickly adapt and adjust recommendation strategies to ensure that recommendation results remain consistent with the user’s current interests and needs. By applying FM during the client update phase, a higher level of personalized recommendation can be achieved.
However, it is important to note that when FM is applied to data that differs significantly from the training distribution, performance degradation can occur. This problem, known as out-of-distribution generalization [93], is a challenge that FM must overcome. In addition, if there are biases in the training data set, FM may learn and reinforce these biases, leading to inequitable results across different data distributions [95, 96, 97]. Although fine-tuning FM requires significantly less resources than training from scratch, it still typically requires significant computational resources for effective fine-tuning and updating [18, 36]. This requirement may limit its use on resource-constrained clients, especially in FRS where each user represents a client, potentially limiting the use of FM due to limited computational resources.
@@ IV-B Communication
In FRS, communication between clients and the server is a critical phase. The primary task at this stage is to transmit model update information from each client to the central server while ensuring data security and communication efficiency during the transmission process. In traditional FRS frameworks, communication between clients and the server typically involves transmitting large amounts of update information, which may include model parameters, gradients, or other relevant statistical data. To ensure efficient and secure communication, common strategies include data compression [98, 99, 100], encrypted transmission [101, 102], and secure aggregation [103, 104, 105]. However, with the increasing data scale and higher privacy protection requirements, these traditional methods face significant challenges [36].
FMs, especially LLMs, possess strong representation capabilities, enabling them to extract key features from model updates through efficient data representation [55, 106] and semantic compression techniques [107, 108]. By removing redundant information and unnecessary data, and retaining only essential information, FM can transform this into a more compact representation, significantly reducing the data transmission volume between clients and the server. Optimizing communication protocols [109] and enhancing data compression and security during this process can markedly improve the overall performance of the system.
@@ IV-C Global Aggregation
In FRS, the global aggregation phase of the server is responsible for combining model updates from all clients into a single global model. This phase is of critical importance in ensuring the overall performance and effectiveness of personalized recommendations. Server global aggregation is characterized by the following key features [110]: First, server global aggregation must handle updates from a large number of clients that contain different data and user behaviors. To ensure the overall quality of the model and the accuracy of the recommendations, the server must effectively integrate this diverse data. Second, the global aggregation process must be sufficiently robust to handle potentially noisy data and anomalous updates. Finally, the scalability of the server’s global aggregation is critical, especially when dealing with a large number of clients and massive amounts of data, requiring the system to efficiently process and aggregate updates. In each iteration, given resource constraints and privacy concerns, the central server may only communicate with a subset of clients. Researches [111, 112] indicate that clients should not participate in successive training rounds, as attackers could potentially infer information about the client from transmitted gradients or model parameters. The client selection strategy determines which clients will participate in the current training round based on factors such as computational power, network stability, data diversity, and data quality [113]. For example, client selection can also be done by random sampling or based on statistical characteristics of the data to ensure that the model learns from diverse data sources, thereby increasing its generalization ability.
Context-aware aggregation represents a critical application of FM in global server aggregation. Traditional aggregation methods, such as naive weighted averaging proposed in FedAvg [3], may not fully exploit the contextual information present in the data. FM, however, is capable of intelligently aggregating updates by understanding the semantics and contextual relationships of each participant’s updates [70]. To illustrate, when FM receives customer updates, it analyzes the contextual information of each update, such as user behavior patterns, preferences, and history. This information enables FM to better comprehend the significance and value of each update and to assign distinct weights accordingly [32]. This context-aware aggregation method can more accurately reflect user preferences and behaviors, thereby enhancing the overall performance of the model. Dynamic weighting permits the system to adapt more effectively to users’ personal needs and to provide recommendations that are more aligned with their interests.
With its extensive knowledge base [114], FM can utilize this knowledge to make more informed decisions during the aggregation process. Knowledge-based aggregation strategies can capitalize on FM’s knowledge and reasoning capabilities to enhance the efficacy of aggregation. To illustrate, in the event that certain customers lack sufficient data, FM can leverage pertinent information from its knowledge base to infer relationships and patterns between data, thereby supplementing and improving aggregation. The key concept of Mixture of Experts (MoE) [115] is that different expert models specialize in different parts of a problem. By appropriately weighting and combining their outputs, better overall performance can be achieved. The MoE architecture has a similar function to aggregation strategies in FRS. Therefore, applying this architecture to the aggregation strategy in FRS may result in improved recommendation performance.
@@
In terms of anomaly detection, FM also performs exceptionally well [116, 117]. In FRS, client updates may be influenced by various factors, leading to noisy or abnormal data [29]. If these abnormal updates are directly used in global aggregation, they may affect the overall quality of the model. FM can utilize its strong anomaly detection capabilities to analyze client updates, identify and handle these anomalies, ensuring the robustness of global aggregation. For instance, if a client’s data update exhibits a notable divergence from the majority of client updates, it may be regarded as noise or an anomaly. FM has the capacity to elect to disregard or assign a diminished weight to it, thus reducing its deleterious impact on the global model and ensuring the stability and robustness of the global model. Liu et al. [37] observed in their study that FM is capable of effectively filtering out noisy data by identifying aberrant patterns, thereby enhancing the stability and reliability of the aggregation process.
This application direction has considerable potential for future development and is worthy of further research and investigation.
V Challenges
The application of FM to FRS marks the advent of a new frontier in personalized content recommendation. This approach promises to further enhance user experience by employing more powerful models than traditional deep models while protecting user privacy. However, this promising combination comes with many challenges, involving various aspects such as data privacy and security, data sparsity and sample imbalance, model training and optimization, model interpretability and transparency, and so on. It is therefore of the utmost importance to gain a comprehensive understanding of these issues and to develop effective strategies to address them. This is essential if FRS is to be successfully deployed and operated in a secure yet powerful manner. In this section, we will examine these challenges in detail, explore their implications and discuss potential strategies to overcome them.
V-A Data-Related Challenges
  • Data Privacy: In a federated setting, the protection of user privacy is of paramount importance. The model must ensure that training and recommendations are performed without disclosing personal data. However, FMs, such as GPT-3, have the potential to memorize and reproduce data from the training set, thereby potentially leaking sensitive information. Additionally, if the generated data is too similar to the original data, there is a risk of privacy infringement. This may involve the use of additional encryption techniques to safeguard user data. Differential privacy [118] protects individual data by adding noise during model training, thus reducing the risk of data leakage while maintaining model performance. Homomorphic encryption [119] allows computations on encrypted data, enabling model training without decrypting the data, effectively preventing data theft during transmission and processing. Furthermore, machine unlearning techniques [120] can remove specific user data from the foundational model, ensuring that these data no longer influence the model’s predictions and outputs. This approach is particularly effective in the context of user deletion requests or in the context of compliance with privacy regulations such as the GDPR.
  • Data Security: The integration of FRS and FM also necessitates the equal consideration of data security. Participants may be subjected to malicious attacks, including Member Inference Attacks [121], Data Reconstruction Attacks [122], and Poisoning Attacks [123], which could compromise the model or alter the data. Consequently, ensuring the integrity and confidentiality of data during transmission and storage represents a significant challenge. The application of secure multi-party computation (MPC) [124] can distribute computational tasks among multiple participants, ensuring that no single party can obtain complete data, thereby enhancing security. Additionally, blockchain [125, 126] can be used to record data access and operation logs, ensuring data integrity and traceability. This also helps improve the transparency and security of the system.
  • Data Sparsity and Scarsity: RS frequently encounters the challenge of data sparsity, where there is limited interaction data between users and items. This leads to decreased model performance. This issue is particularly pronounced in a federated setting, where each participant has limited data. In specific domains or tasks, insufficient data can cause severe data scarcity, further restricting model training and performance. Although FM has the ability to perform zero-shot and few-shot learning, they still need optimization to handle extremely sparse data. It is therefore of great importance to utilise data augmentation techniques [127, 128] in order to generate a greater quantity and variety of training samples, thereby increasing the overall data set. Alternatively, knowledge transfer [129] between different domains can be employed in order to leverage the rich data sets from other areas, thus alleviating the issue of data sparsity and improving the overall model performance in the current domain. Despite the ability of generative FM [130, 131] to produce a vast quantity of synthetic data, the quality of this data may not be guaranteed. If the FM utilises low-quality or biased data sources, the generated data may inherit these issues and even exacerbate existing biases and errors [132]. Wang et al. [22] found that generated data is not always beneficial for representation learning. For example, in contrastive learning, the optimal performance was achieved when the mix ratio of real to generated data was approximately 10:1. Consequently, when generating synthetic data, it is of paramount importance to establish quality standards and utilise evaluation metrics to assess data quality. Furthermore, it is essential to implement quality control mechanisms to filter out low-quality or biased data.
  • Sample Imbalance: In a federated setting, there are significant differences in data size and distribution among different clients [133, 134, 135]. Similarly, in RS, item labels and user behaviors often exhibit a long-tail distribution [136]. These issues are referred to as sample imbalance, where part of classes have much more samples than others in the dataset. This can cause the model to overfit the majority classes while having poor generalization for minority classes, thus affecting the model’s training effectiveness. To address this, additional methods are required to balance these differences and ensure the model’s fairness and accuracy. Resampling techniques typically involve either oversampling or undersampling to balance the dataset [137]. Oversampling increases the number of minority class samples, while undersampling reduces the number of majority class samples. Weighted loss functions can assign higher weights to minority class samples during model training, enhancing the model’s ability to recognize minority classes. This approach can enhance the fairness and accuracy of the model without modifying the data distribution.
V-B Model-Related Challenges
  • Model Synchronization: In a federated setting, models from various participants frequently communicate with the server for synchronization, resulting in high communication costs and complex system management issues. This is particularly problematic for large FM due to their vast number of parameters [18], which require significant computational resources and time for transmission and synchronization. The use of gradient compression techniques [138] to reduce the amount of data transmitted during model synchronization can effectively lower communication costs and enhance synchronization efficiency. Furthermore, the implementation of asynchronous update strategies [139], which permit participants to independently update their models and periodically perform global synchronization, can reduce waiting times during synchronization and enhance the overall efficiency of the system.
  • Model Heterogenity: Model heterogeneity [140] is typically attributable to a multitude of factors, including disparate types, sizes, and architectures of models among various clients. This can be attributed to differences in client hardware, data distribution, or business needs. In certain instances, different clients may utilize distinct model structures. For instance, some clients may employ simple linear models, while others may utilize complex deep learning models [141]. This renders it challenging to deploy or fine-tune a unified foundation model in a federated setting. The design of adaptive training algorithms [142, 143] enables the dynamic adjustment of model parameters based on the computational capabilities and data characteristics of each participant, thereby enhancing the model’s adaptability and performance in heterogeneous environments. Furthermore, knowledge transfer [47] among different participants facilitates the sharing of insights and the mitigation of data and computational resource disparities, resulting in an improvement in the overall performance of the model.
  • Model Interpretability: FM is often represented by black-box models, which makes it challenging to comprehend their internal workings. This lack of interpretability can give rise to issues of trust and compliance in practical applications. In FRS, model interpretability is of paramount importance for establishing user trust and meeting regulatory requirements [144]. Therefore, it is vital to employ explainable AI techniques [145] to provide transparent and understandable explanations for the decisions made by FM. For instance, attention mechanisms [146] and feature importance analysis [147] can be utilized to elucidate model decisions. Furthermore, the utilisation of generative FM [148] to generate natural language explanations for recommendations can provide users with the context and rationale behind the model’s decisions, thereby enhancing user trust and acceptance of the model.
  • Model Transparency: The concept of model transparency entails the open and auditable process of model decision-making. In the context of FRS, transparency plays a pivotal role in fostering user trust, necessitating the assurance of model fairness and non-discrimination. This encompasses the monitoring of model training processes, the validation of model updates, and the tracing of generated data. A lack of transparency can diminish the credibility of the model and increase compliance risks. FRS with FM should have transparent training protocols [149] that record and monitor each step of the model training process. This ensures transparency, thereby enhancing system credibility and compliance. Similarly, establishing a detailed logging system [48] to record every model update and operation ensures traceability, helping to identify and correct issues within the model and improve overall system transparency.

### v3

@@ III-B1 Taxonomy of FM
As shown in Fig. 2, similar to FRS, FMs can also be classified based on data type or functionality. Specifically, based on the type of data used during training, FM is mainly divided into followings [10]:
  • Language FM: Trained on textual data to understand and generate language [18, 53], excelling in natural language processing tasks such as machine translation [59], search [60, 61], and sentiment analysis [62]. Typical language models include BERT [18], GPT-3 [52], and LLama [63], which typically use deep transformer architectures that are able to capture complex linguistic regularities and demonstrate excellent performance on multilingual tasks.
  • Vision FM: Focused on understanding image data, these models, such as DINOv2 [64] and SAM [65], are used for object recognition [19, 66, 67], visual reasoning [68, 69] and image segmentation [65, 70, 71].
  • Multi-modal FM: These models, e.g. CLIP [20] and DALL·E [21], integrate multiple types of data, such as text and images, allowing them to handle cross-modal tasks like image captioning and visual quizzing [72, 73, 74] by integrating information from different modalities.
Functionally, FMs can be categorized into [10, 75, 35]:
  • Discriminative FM: Primarily used for tasks like classification [76, 77] and regression [78], focusing on distinguishing between different classes of input data.
  • Generative FM: Such as GPT-3 and DALL·E, are designed to generate new data instances by learning the distribution of the training data [79], supporting applications like text [80] and image generation [14, 15].
Each type of FM offers unique advantages in its respective area. Language, vision, and multi-modal models provide specialized capabilities for processing different data types, while discriminative and generative models offer distinct approaches for solving problems based on either classification or data generation. By leveraging these specialized models, FMs can address a wide range of challenges across diverse domains.
@@ III-B2 Adaptation
Fig. 1 illustrates a commonly used technique for pre-trained models, known as the Adaptation. Adaptation is a commonly used technique for pre-trained models, where new lightweight layers are added while maintaining the original parameters of the pre-trained model. This enables fine-tuning and extension for specific tasks [81, 38], making it suitable for multi-task learning and tasks performed in resource-limited environments, such as few-shot learning [82]. Adaptation techniques have found wide applications in natural language processing and computer vision tasks [83, 84, 85].
Adaptations can be categorized into three main types [82]:
  • Prompt-based Fine-tuning Adaptation Methods: These methods [86, 87] use prompt learning to adapt models for specific tasks by fine-tuning a portion of the fixed text input as a learnable vector. This allows the model to adjust quickly for new tasks with minimal changes.
  • Adapter-based Fine-tuning Adaptation Methods: Small, trainable modules, called Adapters, are added to a pre-trained model. These adapters [88, 85, 83] are fine-tuned instead of the entire model, making task adaptation efficient with minimal parameter updates.
  • External Knowledge-based Adaptation Methods: External knowledge, such as information from knowledge graphs or additional model queries, is integrated to improve the model’s performance on domain-specific tasks, especially when data is limited [89, 90].
In summary, adaptations play a crucial role in enhancing the versatility of foundation models. By enabling efficient fine-tuning and leveraging external knowledge, these methods help foundation models effectively tackle a wide range of tasks, even in resource-constrained settings or when training data is scarce. This adaptability is key to making foundation models more practical and powerful for real-world applications.
@@ IV Federated Recommendation Systems with Foundation Models
Figure 6: Integration framework of FRSs with FMs, illustrating the key challenges in three stages: Client Model Update, Communication, and Global Aggregation.
As illustrated in Fig. 6, a typical FRS comprises three key stages: client model update, communication, and global aggregation [24]. The integration of FMs aligns with these stages. This section provides a comprehensive overview by summarizing representative works at each stage and examining additional dimensions of integrating FMs into FRS.
@@ IV-A Client Model Update
  • Data Sparsity: As mentioned earlier, a major advantage of FRS is privacy. The data required for model training in FRS, such as user information and interaction history, is privacy sensitive and therefore must remain on the user’s device, creating data silos. This approach not only complies with data privacy regulations (such as GDPR), but also increases user trust. However, this means that each user has a relatively small amount of data compared to the entire dataset, and each user only accesses a portion of the item set, leading to data sparsity.
@@
  • Non-IID Data: Each client independently trains the model on its local data. This allows the model to better adapt to each user’s unique preferences and behaviors, improving personalization and accuracy of recommendations. The resulting data distribution often does not satisfy the independent and identically distributed (IID) assumption, leading to the problem of data heterogeneity.
@@
  • Device Limitations: In addition, user devices are typically consumer products such as mobile phones and personal computers, which are characterized by unstable communication and limited computational resources. This requires keeping the computational load of the models deployed on the client and the amount of information exchanged with the server as low as possible.
@@
FMs are typically pre-trained on large, diverse datasets, providing them with rich knowledge and strong representational capabilities. This pre-training endows them with prior knowledge that can be quickly adapted to specific client data through fine-tuning [10]. Therefore, FMs learn general feature representations, offering adaptability to different data distributions. Applying FMs during the client update phase in FRS can significantly enhance the performance of local models. Specifically, the advantages are outlined as follows:
@@
  • Addressing Data Sparsity: FMs can apply the knowledge learned from large corpora to local data through transfer learning [91]. For instance, a pre-trained FM can be effectively fine-tuned on a small amount of local data at the client to adapt to specific user behaviors, thereby achieving good performance in downstream tasks such as providing more accurate recommendations [35, 92]. Furthermore, by fine-tuning the base model locally, sensitive data does not need to leave the device, protecting user privacy [93]. Thus, even with limited data, clients can achieve better model performance, alleviating the challenges brought by data sparsity.
@@
  • Handling Non-IID Data: FMs’ strong representational capabilities enable them to capture complex user preferences and behaviors. For example, FMs possess powerful semantic understanding abilities, allowing them to better interpret user search queries [60, 61], comments [34], and other textual data [94], thereby providing recommendations that align more closely with user needs. Moreover, FMs can learn complex patterns from non-IID data [11, 95, 96], enhancing the model’s adaptability to different user preferences and ensuring that recommendation results stay consistent with the user’s current interests and needs. Applying FMs during the client update phase can achieve a higher level of personalized recommendation, effectively addressing data heterogeneity.
  • Adapting to Device Limitations: Although fine-tuning and updating FMs typically require significant computational resources, optimizing the model architecture and employing lightweight fine-tuning techniques can reduce the computational burden on devices [97]. The knowledge transfer capability of FMs allows for efficient local fine-tuning under limited computational resources, thereby achieving good recommendation performance even in resource-constrained environments [98, 97]. Additionally, model compression techniques can be used to reduce the model size, thereby lowering computational and storage requirements to suit the limitations of user devices [99].
These characteristics make FMs highly suitable for addressing the diverse data requirements of clients in FRS, thereby improving the generalization performance of local models.
@@ IV-B Communication
Communication between clients and the server is a critical phase in FRSs, involving the transmission of model update information from each client to the central server. Ensuring efficiency and security during this process is essential for maintaining system performance and protecting user privacy. However, this phase faces significant challenges, including the high communication costs associated with transmitting large volumes of data and the inherent security risks of handling sensitive information. Below, we firstly outline these challenges as follows:
  • Communication Costs: The transmission of large-scale model parameters, gradients, or other statistical data can result in excessive communication overhead. This is especially problematic in systems with many clients or high-dimensional models, where bandwidth and latency constraints become bottlenecks.
  • Transmission Security: Sensitive data encapsulated in model updates may be vulnerable to interception during transmission. Such breaches could lead to privacy violations or exposure of user data, undermining the privacy-preserving goals of FRS. Although encryption techniques can mitigate these risks, they often increase the computational and communication burden.
Addressing these challenges requires innovative approaches that enhance both efficiency and security. FMs, with their advanced representation and compression capabilities, offer promising solutions to overcome these limitations:
  • Reducing Communication Costs: FMs can perform semantic compression [100, 101], extracting critical features while eliminating redundant information, thus significantly reducing the data transmitted between clients and the server. Furthermore, optimized communication protocols tailored to FM representations [102] can further improve data exchange efficiency.
  • Enhancing Transmission Security: Transmitting sensitive data in FRSs poses significant privacy risks. FMs enhance transmission security by transforming raw model updates into abstract representations that retain essential information while removing sensitive details, reducing the risk of privacy leakage [33]. Additionally, FMs enable semantic compression [103], which not only reduces communication overhead but also minimizes the risk of adversarial attacks. When combined with techniques such as secure aggregation [104], FMs provide an efficient and scalable solution that strengthens both the security and performance of data transmission in FRS.
@@ IV-C Global Aggregation
In FRS, the global aggregation phase of the server is responsible for combining model updates from all clients into a single global model. This phase is of critical importance in ensuring the overall performance and effectiveness of personalized recommendations. Server global aggregation is characterized by the following key features [105]:
  • Diverse Client Updates: The server must handle updates from a large number of clients, each containing different data and user behaviors. To ensure the overall quality of the model and the accuracy of recommendations, the server must effectively integrate this diverse data.
  • Noisy Data: The global aggregation process must be sufficiently robust to handle potentially noisy data and anomalous updates from clients, which can adversely affect the global model’s performance.
  • Scalability: The scalability of the server’s global aggregation is critical, especially when dealing with a large number of clients and massive amounts of data. The system must efficiently process and aggregate updates to maintain performance.
@@
To address these challenges, several existing strategies have been developed. For example, in each iteration, given resource constraints and privacy concerns, the central server may only communicate with a subset of clients. Studies [106, 107] indicate that clients should not participate in successive training rounds, as attackers could potentially infer information about the client from transmitted gradients or model parameters. The client selection strategy determines which clients will participate in the current training round based on factors such as computational power, network stability, data diversity, and data quality [108]. For example, client selection can be done by random sampling or based on statistical characteristics of the data to ensure that the model learns from diverse data sources, thereby increasing its generalization ability. However, these existing strategies have limitations, such as suboptimal aggregation efficiency and challenges in fully exploiting the contextual information present in client updates.
FMs can play a significant role in enhancing the global aggregation process. Traditional aggregation methods, such as naive weighted averaging proposed in FedAvg [3], may not fully exploit the contextual information present in the data. FMs, however, are capable of intelligently aggregating updates by understanding the semantics and contextual relationships of each participant’s updates [70].
  • Context-Aware Aggregation: FMs utilize their powerful representation capabilities to analyze the structural and statistical characteristics of client updates, including update frequency, scale, and trends. This enables the server to more effectively assess the significance of each update without relying on potentially sensitive user-specific information, thereby aligning the aggregation process more closely with system-wide patterns and requirements [29]. By understanding the latent relationships between updates, FMs assign appropriate weights to individual contributions, improving the fidelity of the global model.
  • Dynamic Weighting: FMs enable dynamic adjustment of update weights based on their relevance to the global model [97]. For instance, updates contributing to under-represented data distributions may receive higher weights, ensuring that the global model remains inclusive and robust to diverse data sources [109].
  • Knowledge-Based Aggregation: Leveraging the extensive knowledge base of FMs [110], knowledge-based aggregation strategies have the ability to capitalize on FMs’ knowledge and reasoning capabilities to enhance the aggregation process. For instance, if certain clients lack sufficient data, FMs can use pertinent information from their knowledge base to infer relationships and patterns, thereby supplementing and improving the aggregation process to have better performance [111].
@@
  • Anomaly Detection and Handling: FMs also excel in anomaly detection [112, 113]. Client updates may be influenced by various factors, leading to noisy or abnormal data [26]. If these abnormal updates are directly used in global aggregation, they may affect the overall model quality. FMs can analyze client updates to identify and handle anomalies, ensuring the robustness of global aggregation. For example, if a client’s data update shows a notable divergence from the majority, it may be regarded as noise or an anomaly, and FMs can choose to disregard or assign a reduced weight to it, thereby maintaining the stability of the global model [34].
Overall, FMs effectively tackle the key challenges in the global aggregation phase by integrating client updates with context-awareness, dynamically weighting to handle noisy data, leveraging pre-trained knowledge for scalability, and performing robust anomaly detection. These approaches significantly enhance the global model’s quality, robustness, and adaptability, positioning FMs as a promising tool for advancing FRSs. Continued research in this area could further unlock their potential, fostering progress in personalized, scalable, and privacy-preserving recommendations.
V Key Challenges and Strategic Solutions
Figure 7: Challenges in integrating FMs into FRSs, divided into data and model challenges, highlighting areas requiring strategic solutions for enhanced personalization and privacy.
Integrating FMs into FRSs has introduced new opportunities for enhancing personalized recommendations while maintaining user privacy. However, as illustrated in Fig. 7, this integration presents significant challenges across multiple dimensions. Understanding these challenges comprehensively and developing effective strategies to address them is crucial for advancing FRSs. This section explores these challenges in depth and presents potential strategies to overcome them.
V-A Data Challenges and Mitigation Strategies
V-A1 Privacy Concerns
Privacy aims to prevent unauthorized access or misuse of data, especially during model training and data handling, to protect user identities and personal details [114]. In a federated setting, safeguarding user privacy is of utmost importance. Privacy concerns focus on preventing unintended exposure or misuse of personal data, ensuring that sensitive information remains confidential. While training and recommendations must occur without revealing personal data, FMs, such as GPT-3, may memorize and reproduce training data, potentially leaking sensitive information [115]. Additionally, if generated data closely resembles original data, user privacy risks may arise [114].
To mitigate these privacy concerns, differential privacy [116] can be employed by adding noise during model training to reduce data leakage risks while preserving model performance. Homomorphic encryption [117] allows training on encrypted data, effectively preventing data theft during transmission and processing. Moreover, machine unlearning techniques [118] can be utilized to remove specific user data from the FMs, ensuring compliance with privacy regulations like GDPR. An emerging approach, known as privacy rewrite, transforms sensitive data into anonymized formats before transmission. Frameworks like HaS [115] exemplify this by anonymizing private entities locally and reconstructing them after processing. This method reduces privacy risks, preserves data utility, and minimizes computational overhead.
V-A2 Security Threats
Security threats involve protecting data from malicious attacks or breaches [114]. While privacy controls who can access and use data, security emphasizes maintaining data integrity and defending against threats such as cyberattacks [119]. The integration of FMs into FRSs also necessitates ensuring data security, as participants may be subject to attacks like Member Inference Attacks [119], Data Reconstruction Attacks [120], and Poisoning Attacks [121], which can compromise model integrity or data quality.
To ensure data integrity and confidentiality during transmission, secure multi-party computation [122] can distribute computational tasks among multiple participants, preventing any single party from accessing complete data. Additionally, blockchain [123, 124] can record data access and operations, enhancing data integrity, traceability, and transparency.
V-A3 Data Scarcity
Despite FMs’ potential to alleviate data sparsity, FRSs often face data scarcity, especially in scenarios where participants have limited interaction data under federated Settings [24]. Data sparsity can significantly impact model performance, particularly when user interactions are insufficient to build effective models.
Data augmentation techniques [125, 126] can be used to generate diverse training samples, expanding the dataset size. Knowledge transfer [98] from rich datasets in other domains can also alleviate data sparsity, improving overall model performance. When employing generative FMs [127, 128], quality control mechanisms are essential to ensure that synthetic data meets quality standards without introducing biases.
V-A4 Data Imbalance
In federated settings, significant differences in data size and distribution exist across clients, causing data imbalance [109, 129, 130]. This imbalance, often driven by long-tail distributions of item labels and user behaviors, affects model training effectiveness.
To address sample imbalance, resampling techniques like oversampling minority classes or undersampling majority classes [131] can be used. Weighted loss functions [132, 109] can also be used to assign higher importance to minority class samples during training, improving the capability to recognize underrepresented data without altering the data distribution.
V-B Model Challenges and Strategic Solutions
V-B1 Synchronization Overhead
In federated settings, the need for frequent synchronization between participants and the central server can result in high communication costs and system complexity, especially when dealing with large FMs with extensive parameters [10].
Gradient compression [133] can help reduce the data transmitted during synchronization, thereby lowering communication costs. Asynchronous update strategies [134] that allow participants to update models independently, followed by periodic global synchronization, also enhance system efficiency.
V-B2 Client Heterogeneity
Arising from variations in model types, sizes, and architectures, this issue creates challenges when deploying a unified FM in federated settings [135].
Adaptive training algorithms [136, 137] that adjust model parameters based on each client’s computational capabilities and data characteristics can address this issue. Knowledge transfer [22] between clients can further enhance model performance by mitigating disparities.
V-B3 Interpretability and Transparency
FMs are often perceived as black-box models, which makes it challenging to understand their internal workings and may lead to issues with trust and compliance [33].
Explainable AI [138], such as attention mechanisms [139] and feature importance analysis [140], offer insights into model decisions, enhancing transparency. Additionally, generative FMs [141] generate natural language explanations for recommendations, fostering user trust and acceptance.
Figure 8: Key future directions for integrating FMs into FRSs aim to address emerging challenges and opportunities, focusing on enhancing recommendation performance, safeguarding user privacy, and improving system adaptability.